# Model for the Thermodynamics of Iron at High Pressures Near Melting

### v1

@@ Abstract
The Fe pressure-temperature phase diagram and its melting line have a wide range of applications, including providing constraints for iron-core planetary models. We propose an equation of state (EOS) model based on the interstitial theory of simple condensed matter (ITCM) suggested by A.V. Granato [1]. When applied to Fe, this model allows for the extrapolation of measured melting lines to Earth’s inner core boundary (ICB) conditions. The ITCM describes the solid-liquid phase transition in metals as resulting from a strong structural perturbation due to a high concentration of interstitial-like defects. The strong nonlinearity of their self-interaction causes the stabilization of this interstitial-rich phase. The original model is expanded to describe melting over a wide range of pressures and temperatures rather than focusing on a specific isobaric transition. Using this model, we fit the measured melting data, extrapolate it to cover ICB conditions, and develop a multiphase equation of state (EOS) that encompasses this regime. The model is used to explain contradictory data regarding the location of the melting line, resulting from a novel phase transition between two separate liquid phases, specifically between FCC-based and HCP-based liquids. This additional liquid phase offers a new interpretation of the previously suggested near-melting high-pressure phase [2] and may also provide a solution to the inner core nucleation paradox [3].
@@
Introduction. — Iron’s behavior at extreme pressures and temperatures is important in materials science and geophysics. Iron, the most abundant element in Earth’s core, has a melting line and thermodynamic properties essential to understanding processes such as planetary formation, magnetic field generation, and seismic activity [4, 5, 6, 7, 8, 9, 10].
The melting temperature of iron under these conditions is evaluated using dynamic shock loading [11, 12, 13, 14] and laser-heated diamond anvil cells (DAC) [15, 16]. While these techniques enable access to high temperatures and pressures, they yield significantly different results, with discrepancies of up to 3000 K in the estimated Fe melting temperature at Earth’s inner core boundary (ICB) conditions [17]. These variations primarily arise from experimental limitations, such as limited access to combined high temperatures and pressures in DAC measurements and kinetic effects in shock wave experiments that may lead to observations of metastable states [18, 19]. The extrapolation of the melting line relies on fitting models to theoretical predictions, including density functional theory (DFT) calculations [20, 21] and multiphase equation of state calculations [22, 7, 23]. Thus models for the equation of state (EOS) of iron are needed both to extrapolate from experimental and theoretical specific evaluations as well as a platform to identify and explain contradictions between such values.
@@
In this manuscript, we develop an EOS model to address these discrepancies and enhance our understanding of iron under ICB conditions. A significant limitation in developing such a model is the scarce availability of experimental data in the warm, dense regime. Our aim is to construct a thermodynamically consistent model for Fe at the ICB conditions by extrapolating from available measurements of lower-temperature dense solid phases to higher temperatures. To that end we base our work on a class of models [1, 24, 25, 26, 27, 28, 29] that describe the free energy difference between the solid and liquid phases by treating the liquid as a highly perturbed solid and estimation explicitly the variation of the free energy due to a specific microscopic perturbation. Consequently, these methods may allow for improved extrapolation of phase lines into regions that are not directly measured.
While a wide range of possible perturbations were considered in the past, many of these suffer from different contradictions, such as too high energy per atomic length energy to be generated thermally for dislocations [29] or too low entropy per volume in vacancy-based models[28].
We employed an EOS model based on Granato’s interstitial theory of simple condensed matter (ITCM) [30, 1, 31]. This theory explains isobaric melting as a transition to a state with a high concentration of interstitial-like defects. The stabilization of the liquid, defect-rich phase is due to the strong nonlinearity of their self-interaction. Furthermore, Granato’s model [1] was shown to be consistent with transport and kinetic processes. For example, solids with high interstitial concentrations simulated using Molecular Dynamics (MD) showed similar dynamics to those of liquids[32]. In addition, diffusion of interstitial-like defects was shown to control the crystallization velocity[33] with simulation being consistent with experiments at large supercooling [34].
@@
Free Energy Model. — We extend the original model [31, 1, 30], which describes an isobaric phase transition, to a generalized form that covers a range of pressures and reconstructs the melting line over a broad thermodynamic regime.
@@
where $V$, $P$ and $T$ are the volume, pressure and temperature, respectively, $B$ and $\mu$ are the bulk and shear moduli and $\alpha_{1}$ and $\alpha_{2}$ are linear coefficients, it is assumed that $\alpha_{1}\gg\alpha_{2}$. This is combined with a model, that most changes in the free energy due to strain are caused by changes in the formation enthalpy, which leads to the following expression for the shear modulus: $\mu(P,T,c)=\mu_{0}(P,T)\exp(-\alpha_{1}\beta c)$, Where $\beta$ is a dimensionless free parameter that depends on the crystal structure, the resulting softening of the shear modulus with the intersitial concentration is consistent with experimental observations. [1]
The expression in Eq. 2 depends on the volume, for which we assume an exponential dependence on the interstitial concentration: $V(P,T,c)=V_{0}(P,T)\cdot e^{\xi c}$. Where $\xi$ is a free parameter quantifying the relationship between interstitial concentration and volume change. This approximation was verified to produce a consistent expression with the volume obtained from the derivative of the free energy.
@@
Fitting. — The Helmholtz free energy for each solid phase is written as a combination of three terms:
@@
Where $F_{cold}$ represents the $T$-independent ”cold energy” and $F_{ion},F_{electron}$, which represent changes to the free energy from ionic and electronic excitations, respectively. Treating ionic and electronic excitations as independent is a manifestation of the Born–Oppenheimer approximation, as is often assumed in EOS studies[35, 36, 22, 37].
@@
where $C_{V},\beta_{e}\text{ and }\gamma_{e}$ are the electronic heat capacity, electronic heat capacity coefficient and the electronic Grüneisen coefficient, respectively.
@@
We fitted the free energies for the $\alpha$ (BCC), $\gamma$ (FCC) and $\varepsilon$ (HCP) phases to various experimental data. This includes: 1-bar T-dependent thermal-expansion [40, 41] and heat capacity (entropy) [42] data for the $\alpha$ and $\gamma$ phases, Diamond Anvil Cell (DAC)[43] and principal Hugoniot measurements [44, 45], for $\alpha$ and $\varepsilon$ phases, heated Hugoniot [46] for $\gamma$, ramp-compression data [47] for $\varepsilon$ and the phase transition lines data [48, 49]. The fitted parameters are listed in [50]
@@
To model the liquid state EOS, we use the fitted multiphase solid EOS and the change in free energy due to the high concentrations of interstitials in Eq. Model for the Thermodynamics of Iron at High Pressures Near Melting. Since the free energy of the liquid is described as a perturbed crystalline state, the fitting of the free energy was performed by identifying a solid state near melting and fitting a liquid phase transition using this phase as a reference. The properties of the underlying crystalline phase are predetermined by the fit within its thermodynamic stability range. This leaves as free-fitting parameters only those that describe interactions between the interstitials at high concentrations. We fitted liquid phases using only for $\gamma$ (FCC) and $\varepsilon$ (HCP) melting lines, which dominate at high pressures and have been previously fitted. The BCC-related $\alpha$ and $\delta$ phases are stable at lower pressures and are excluded due to their limited relevance. This fitting leads to two separate liquid free energies, namely, FCC- and HCP-based liquids, denoted as $L_{FCC}$ and $L_{HCP}$.
@@
The model was fitted to the melting line measurements from Sinmyo [16], as this dataset offers the broadest pressure range and closely approximates ICB conditions, enabling the most reliable extrapolation of the melting line. To maintain thermodynamic consistency, Richard’s rule [51] was incorporated, governing the entropy change during melting. Additionally, bulk and shear moduli were independently fitted to a second-order pressure and first-order temperature expressions based on ab initio calculations [52, 53, 54], allowing for extrapolation beyond the regime where measurements of the solid phases are available.
@@
Figure 1: Fe phase diagram derived from our multiphase EOS. Light green indicates the $\alpha(BCC)$ phase, blue indicates $\gamma(FCC)$, red indicates $\varepsilon(HCP)$, light blue indicates $L_{FCC}$, pink indicates $L_{HCP}$. White markers indicate the Shock melt experiments where each marker indicates a different experiment: square[11], diamond[12], circle[13] and triangle[14]. The black dashed line indicates the $\gamma-\varepsilon$ phase transition[49]. The orange[16] and green[15] circle indicates the DAC melt experiments.
@@
Results and discussion. — A comparison between the phase diagram resulting from the fitted EOS to other calculations and measurements of the Fe melting line is presented in Fig. 1. The model was fitted to and reproduces melting data from DAC measurements by Sinymo[16] and disagrees with previous measurements by Anzellini [15]. Extrapolating the liquid phase free energies expression beyond their fitting range produced a hypothetical $L_{FCC}-L_{HCP}$ phase transition between the two liquid phases. The existence and location of this transition in the P-T phase diagram are not sensitive to a specific choice of fitting parameters. This phase transition agrees with results from shock melting experiments [11, 14, 13], and may explain the discrepancy between melting measurements at high pressure. Specifically, it reproduces the explicit observation of two-phase transitions measured by Brown& McQueen [11]. Such first-order liquid-liquid phase transitions were previously observed in experiments and atomistic simulations of other liquids, including liquid metals[55, 56].
@@
Figure 3: The Hugoniot possible trajectories near melting. The red dotted lines indicate the $\varepsilon-L_{HCP}$ and the $L_{HCP}-L_{FCC}$ phase transitions. The solid blue line represents the derived Hugoniot trajectory, while the dashed blue line indicates a metastable branch that passes through the edges of the solid stability regime. The markers represent different shock melting measurements: gray circles[11], light blue diamonds[12], green triangles[13] and yellow squares[14].
@@
In Fig. 3, the Hugoniot curves derived from the model are compared with shock melt measurements. The EOS leads to two closely spaced phase transition lines near melting at high pressures, one for melting and the other for the liquid-liquid phase transition. These two phase transitions are expected to lead to two separate transitions in the measured Hugoniot trajectory as it passes through the transition zone. Such a double transition was reported in shock compression measurement of elastic velocities [11]. The theoretical Hugoniot curve passing through these phases is shown in Fig. 3, which agrees with measurements below and above the melting and the liquid-liquid phase transition lines. Between these lines, kinetics is expected to play a significant role as persistent metastable states are expected to be observed in dynamic experiments. The dashed line in Fig. 3 represents the Hugoniot trajectory continuing in metastable solid state until the upper limit of solid metastability. This leads to an improved fit to the measured data at $250(GPa)$[11, 13, 14]. The optional trajectories corresponding to different transition points may explain the spread in the experimental data. Specifically, depending on kinetics, even a single transition from metastable solid to $L_{FCC}$, completely bypassing the intermediate liquid state, is possible. The measurements by Yoo and Holmes [12] deviate from our EOS, as shown in Figs. 1 and 3. This discrepancy may be attributed to the model used in [12] for the thermal capacity of the solid state, leading to differences in the estimated temperature, which is not directly measured.
@@
In addition, seismic velocity measurements and simulations have shown that Earth’s inner core displays strong shear softening and an ultrahigh Poisson’s ratio. In simulations of Fe HCP solid close to melting, this softening was attributed to the formation of increased correlation in the displacement of atoms, manifested in chains of collective motion.[57] Such increased correlation in displacement chains was previously reported in atomistic simulations of Cu’s highly perturbed crystalline phase as melting was approached.[32] Thus, this observed softening may be related to the suggested intermediate $L_{HCP}$. The formation of such an intermediate liquid phase is expected to reduce the nucleation barrier due to its effect on the solid-liquid interface energy[33, 34]. This, in turn, can explain the gap between the observed structure and the estimated nucleation time of the solid phase in Earth’s inner core. [3]. Furthermore, the $L_{HCP}$ phase is a natural candidate as an intermediate phase for the two-step process, which was previously suggested to explain this phenomenon [58].
@@
Another type of thermodynamic trajectories, that may be accessible experimentally, are the compression isotherms of the EOS presented in Fig. 4. The model’s entropy jump at melting is consistent with Richards’ rule [51], and the density profile follows a reasonable trend. [43]
@@
Conclusions. — Using ITCM, we derived expressions for the free energy functions of various iron phases at high pressure and temperature. Applying ITCM with two different solid reference phases at low temperatures leads to a phase diagram that includes two liquid phases. Calculations based on these functions coincide with measurements and prior computations.

### v2

@@ Abstract
The Fe pressure-temperature phase diagram and its melting line have a wide range of applications, including providing constraints for iron-core planetary models. We propose an equation of state (EOS) model based on the interstitial theory of simple condensed matter (ITCM) suggested by A.V. Granato [1]. When applied to Fe, this model allows for the extrapolation of measured melting lines to Earth’s inner core boundary (ICB) conditions. The ITCM describes the solid-liquid phase transition in metals as resulting from a strong structural perturbation due to a high concentration of interstitial-like defects. The nonlinearity of their self-interaction leads to the stabilization of this interstitial-rich phase. The original model is expanded to describe melting over a wide range of pressures and temperatures rather than focusing on a specific isobaric transition. The model is fitted to measured melting data, and then extrapolated to cover ICB conditions. The suggested application of ITCM naturally leads to the suggestion of two distinct liquid phases, namely HCP- and FCC-based liquids. We show that these phases and the transition between them are consistent with previous thermodynamic measurements. The additional liquid phase offers a new interpretation of the previously suggested near-melting high-pressure phase [2] and may also explain the inner core nucleation paradox [3].
@@ I Introduction
Iron’s behavior at extreme pressures and temperatures is important in materials science and geophysics. Iron, the most abundant element in Earth’s core, has a melting line and thermodynamic properties essential to understanding processes such as planetary formation, magnetic field generation, and seismic activity [4, 5, 6, 7, 8, 9, 10].
Over the past decades, considerable effort has been devoted to constraining the melting temperature of Fe under Earth’s core conditions. Dynamic shock-loading experiments [11, 12, 13, 14, 15] and laser-heated diamond anvil cells (DAC) [16, 17, 18, 19, 20] have provided direct experimental access to extreme pressures and temperatures. However, these methods often yield markedly different results, with discrepancies of up to 3000 K in the estimated melting temperature at the inner core boundary (ICB) [18]. These variations largely reflect methodological limitations: laser-heated DAC experiments have limitations in temperature and pressure measurement [21], while shock experiments may probe metastable states due to kinetic effects [22, 23].
Melting of Iron at high pressures was also studied extensively using Density functional theory (DFT) calculations [24, 25]. The DFT results demonstrate a significant gap, and some of these results were incorporated into the fitting procedure of multiphase EOS models [26, 7, 27]. Possible explanation for the divergence between experimental observations is based on the formation of a stable high-pressure BCC phase, which was suggested theoretically [28, 29] and might be observed in recent experiments. [30] However, the stabilization of such a phase and its possible role in non-pure systems remains unclear [31].
@@
In this manuscript, we pursue a complementary route to explore the properties of Iron at these conditions. We adopt a phenomenological EOS framework, fitting theory to previously available data, assuming there is no additional high-pressure phase. Our work builds on a class of models [1, 32, 33, 34, 35, 36, 37] that describe the free-energy difference between solid and liquid phases by treating the liquid as a highly perturbed solid and by explicitly estimating the free-energy variation associated with microscopic perturbations. Consequently, these methods may allow for improved extrapolation of phase lines into regions that are not directly measured.
A variety of perturbations have been considered in this context. Still, many suffer from inherent contradictions—for instance, dislocation-based models require too high energy per atomic length to be generated thermally [37]. In contrast, vacancy-based models yield entropy per volume that is too low [36]. To overcome these limitations, we focus on Granato’s interstitial theory of condensed matter (ITCM) [38, 1, 39], in which melting is described as a transition to a state with a high concentration of interstitial-like defects stabilized by the strong nonlinearity of their self-interactions. This theory has been shown to reproduce the kinetic and thermodynamic properties of simple liquid metals. Molecular dynamics simulations of solids with high interstitial concentrations reproduce dynamics similar to those of liquids [40], and crystallization speeds at large supercooling were shown to be consistent with interstitial-limited diffusion both theoretically [41] and experimentally [42].
II Free Energy Model
@@
We extend the original model [39, 1, 38], which describes an isobaric phase transition, to a generalized form that covers a range of pressures and reconstructs the melting line over a broad thermodynamic regime. This extension enables the extrapolation of the theory to higher pressures than those where the melt line is reliably known.
@@
where $V$, $P$ and $T$ are the volume, pressure and temperature, respectively, $B$ and $\mu$ are the bulk and shear moduli and $\alpha_{1}$ and $\alpha_{2}$ are linear coefficients, it is assumed that $\alpha_{1}\gg\alpha_{2}$. This is combined with a model, that most changes in the free energy due to strain are caused by changes in the formation enthalpy, which leads to the following expression for the shear modulus: $\mu(P,T,c)=\mu_{0}(P,T)\cdot\exp{(-\alpha_{1}\beta c)}$, Where $\beta$ is a dimensionless free parameter that depends on the crystal structure, the resulting softening of the shear modulus with the intersitial concentration is consistent with experimental observations. [1]
The expression in Eq. 2 depends on the volume, for which we assume an exponential dependence on the interstitial concentration: $V(P,T,c)=V_{0}(P,T)\cdot\exp{(\xi c)}$. Where $\xi$ is a free parameter quantifying the relationship between interstitial concentration and volume change. This approximation was verified to produce a consistent expression with the volume obtained from the derivative of the free energy.
@@
Where $\gamma$ is the Grüneisen coefficient, and $a$, $b$, and $u$ are parameters that depend on the material[1].
@@
III Fitting
III.1 Solid Phases
Developing the EOS and phase diagram requires first fitting the solid free energy functions, which serve as the reference states for the liquid phases. We use a classical and simple analytic form, similar to those shown to be reliable at high pressures and temperatures [26, 7]. Accordingly, the Helmholtz free energy for each solid phase is written as a combination of three terms:
@@ III.1 Solid Phases
Where $F_{cold}$ represents the $T$-independent ”cold energy” and $F_{ion},F_{electron}$, represent changes to the free energy from ionic and electronic excitations, respectively. Treating ionic and electronic excitations as independent is a manifestation of the Born–Oppenheimer approximation, as is often assumed in EOS studies [43, 44, 26, 45].
@@
where $C_{V},\beta_{e}\text{ and }\gamma_{e}$ are the electronic heat capacity, electronic heat capacity coefficient, and the electronic Grüneisen coefficient, respectively.
@@
We fitted the free energies for the $\alpha$ (BCC), $\gamma$ (FCC), and $\varepsilon$ (HCP) phases to various experimental data. This includes: 1-bar T-dependent thermal-expansion [48, 49] and heat capacity (entropy) [50] data for the $\alpha$ and $\gamma$ phases, Diamond Anvil Cell (DAC)[51] and principal Hugoniot measurements [52, 53], for $\alpha$ and $\varepsilon$ phases, heated Hugoniot [54] for $\gamma$, ramp-compression data [55] for $\varepsilon$ and the phase transition lines data [56, 57]. The fitted parameters, fit results, and comparisons are in Appendix. A
III.2 Liquid Phases
@@
To model the liquid state EOS, we use the fitted multiphase solid EOS and the change in free energy due to the high concentrations of interstitials in Eq. II. Since the free energy of the liquid is described as a perturbed crystalline state, the fitting of the free energy was performed by identifying a solid state near melting and fitting a liquid phase transition using this phase as a reference. The properties of the underlying crystalline phase are predetermined by the fit within its thermodynamic stability range. This leaves as free-fitting parameters only those that describe interactions between the interstitials at high concentrations. We fitted liquid phases using only $\gamma$ (FCC) and $\varepsilon$ (HCP) melting lines, which dominate at high pressures and have been previously fitted. The BCC-related $\alpha$ and $\delta$ phases are stable at lower pressures and are excluded due to their limited relevance. The stable phase at each pressure and temperature was determined by identifying the phase with the minimum energy as a function of interstitial concentration for the explored solid phases. This fitting leads to two separate liquid phases, namely, FCC and HCP-based liquids, denoted as $L_{FCC}$ and $L_{HCP}$.
The model was fitted to the melting line measurements from Sinmyo [17], as this dataset offers the broadest pressure range and approximates ICB conditions, enabling the most reliable extrapolation of the melting line. Because these measurements are the only ones above 200(GPa)200~(GPa), they dictate the high-pressure behavior of the fit when additional DAC datasets are included [16, 18, 19]. In practice, adding these datasets produced only negligible changes, as the Sinmyo points lie within their overall spread.
@@
To maintain thermodynamic consistency, Richard’s rule [58] was incorporated, governing the entropy change during melting. Additionally, bulk and shear moduli were independently fitted to a second-order pressure and first-order temperature expressions based on ab initio calculations [59, 60, 61], allowing for extrapolation beyond the regime where measurements of the solid phases are available.
@@
Figure 1: Fe phase diagram derived from our multiphase EOS. Light green indicates the $\alpha(BCC)$ phase, blue indicates $\gamma(FCC)$, red indicates $\varepsilon(HCP)$, light blue indicates $L_{FCC}$, pink indicates $L_{HCP}$. White markers indicate the Shock melt experiments, where each marker indicates a different experiment: square[11], diamond[12], circle[13], downward triangle[14], and upward triangle[15]. The black dashed line indicates the $\gamma-\varepsilon$ phase transition[57]. The orange[17], green[16], light green[18] and gray[19] circle indicates the heated DAC melt experiments. The light blue star indicates heated DAC stability measurements of $\varepsilon(HCP)$[20].
@@
IV Results and discussion
A comparison between the phase diagram derived from the fitted EOS and previous calculations and measurements of the Fe melting line is shown in Fig. 1. The melting line agrees with measurements by Sinmyo [17], while showing some deviation from earlier results by Anzellini, Boehler, and Aquilanti [16, 18, 19]. Overall, it lies near the center of the experimental spread. Moreover, the phase boundaries are consistent with the high-pressure stability of $\varepsilon$-Fe (HCP) measurements reported by Tateno [20].
Extrapolating the liquid phase free energies expression beyond their fitting range produced a hypothetical $L_{FCC}-L_{HCP}$ phase transition between the two liquid phases. The existence and location of this transition in the P-T phase diagram are not sensitive to a specific choice of fitting parameters. This phase transition agrees with results from shock melting experiments [11, 14, 13, 15], and may explain the discrepancy between melting measurements at high pressure. Specifically, it reproduces the explicit observation of two-phase transitions measured by Brown& McQueen [11]. Such first-order liquid-liquid phase transitions were previously observed in experiments and atomistic simulations of other liquids, including liquid metals[62, 63].
@@ IV Results and discussion
Figure 3: The Hugoniot possible trajectories near melting. The red dotted lines indicate the $\varepsilon-L_{HCP}$ and the $L_{HCP}-L_{FCC}$ phase transitions. The solid blue line represents the derived Hugoniot trajectory, while the dashed blue line indicates a metastable branch that passes through the edges of the solid stability regime. The markers represent different shock melting measurements: gray circles[11], light blue diamonds[12], green triangles[13], yellow squares[14], and pink stars[15].
@@
In Fig. 3, the Hugoniot curves derived from the model are compared with shock melt measurements. The EOS leads to two closely spaced phase transition lines near melting at high pressures, one for melting and the other for the liquid-liquid phase transition. These two phase transitions are expected to result in two separate transitions in the measured Hugoniot trajectory as it passes through the transition zone. Such a double transition was reported in shock compression measurement of elastic velocities [11]. The theoretical Hugoniot curve passing through these phases is shown in Fig. 3, which agrees with measurements below and above the melting and the liquid-liquid phase transition lines. Between these lines, kinetics is expected to play a significant role as persistent metastable states are expected to be observed in dynamic experiments. The dashed line in Fig. 3 represents the Hugoniot trajectory continuing in metastable solid state until the upper limit of solid metastability. This leads to an improved fit to the measured data at 250(GPa)250(GPa)[11, 13, 14, 15]. The optional trajectories corresponding to different transition points may explain the spread in the experimental data. Specifically, depending on kinetics, even a single transition from metastable solid to $L_{FCC}$, completely bypassing the intermediate liquid state, is possible. The measurements by Yoo and Holmes [12] deviate from our EOS, as shown in Figs. 1 and 3. This discrepancy may be attributed to the model used in [12] for the thermal capacity of the solid state, leading to differences in the estimated temperature, which is not directly measured.
Additional Shock-Ramp melting line experiments were conducted at the National Ignition Facility (NIF) by Kraus et al. [64]. They determined the phase at each measurement point from the momentum transfer spectra. Subsequent ab initio work [28] observed that a measurement interpreted as a liquid showed deviations from classical liquid theory at both low and high frequencies, leading to an attribution of the data to a BCC phase. Instead, we propose that these deviations could be explained by resonant and local vibrational modes, which are expected within ITCM.
@@
In addition, seismic velocity measurements and simulations have shown that Earth’s inner core displays strong shear softening and an ultrahigh Poisson’s ratio. In simulations of Fe HCP solid close to melting, this softening was attributed to the formation of increased correlation in the displacement of atoms, manifested in chains of collective motion.[65] Such increased correlation in displacement chains was previously reported in atomistic simulations of Cu’s highly perturbed crystalline phase as melting was approached.[40] Thus, this observed softening may be related to the suggested intermediate $L_{HCP}$. The formation of such an intermediate liquid phase is expected to reduce the nucleation barrier due to its effect on the solid-liquid interface energy[41, 42]. This, in turn, can explain the gap between the observed structure and the estimated nucleation time of the solid phase in Earth’s inner core. [3]. Furthermore, the $L_{HCP}$ phase is a natural candidate as an intermediate phase for a two-step process, which was previously suggested to explain this phenomenon [66].
@@
Another type of thermodynamic trajectory that may be accessible experimentally is the compression isotherms of the EOS presented in Fig. 4. The model’s entropy jump at melting is consistent with Richards’ rule [58], and the density profile follows a reasonable trend. [51]
@@
V Conclusions
Using ITCM, we derived expressions for the free energy functions of various iron phases at high pressure and temperature. Applying ITCM with two different solid reference phases at low temperatures leads to a phase diagram that includes two liquid phases. Calculations based on these functions coincide with measurements and prior computations.